# Foundation Cures Personalization: Improving Personalized Models’ Prompt Consistency via Hidden Foundation Knowledge

### v2

@@ 4.2 Asymmetric Prompt Guidance
Following attributes with clear location restored via FASA, to enhance more abstract attributes such as expressions, we introduce a simple yet effective method called asymmetric prompt guidance (APG). This strategy is based on the diffusion model’s inversion  [40]. During the inversion phase, the model accepts only a template prompt (e.g. “a man”) that does not include any target attributes. In the denoising process, we add the target attributes’ tokens back to this template prompt (e.g. “a man” $\rightarrow$ “a man laughing”). By leveraging the pretrained controllability of the foundation model, this approach enhances such attributes, resulting in the final refined image $I_{out}$. Throughout the denoising process, we use only pure textual prompts without identity embeddings, thereby avoiding their potential influence on the tokens related to the target attributes, as discussed in Section 3. Furthermore, to better preserve the identity, we start the denoising process directly from an intermediate latent code $\hat{z_{\gamma T}},\gamma\in[0,1]$, where the high-level identity information has already been established.
@@ 5 Experiments
Figure 6: Qualitative comparison with facial personalization baselines. Different attributes in prompts are highlighted in various colors. Comparison of corresponding FD outputs is provided in the Section 9 of the supplementary material.
Evaluation Datasets. We collect an extensive dataset encompassing both identity and prompt. The identity dataset comprises $50$ identities, with $30$ derived from the CelebA-HQ [25] and the other $20$ non-celebrity identities curated by our team. Each identity is represented by a single image, and the identities’ gender is balanced. Additionally, the identity dataset encompasses a spectrum of facial characteristics which better evaluates the stability of attribute enhancements. The prompt set consists of $20$ prompts: eight incorporate a single attribute, eight incorporate two attributes, and four incorporate three attributes. For each $<$identity, prompt$>$ pair, we produce 20 images for the evaluation process. Detailed information can be found in Section 8 of the supplementary material.
Baselines. We test FreeCure with five representative methods for facial personalization: PhotoMaker [27], PuLID [14], Face-diffuser [45], Face2Diffusion  [39] and FastComposer [47]. PuLID and PhotoMaker use SD-XL as their main foundation models, whereas Face-diffuser, Face2Diffusion, and FastComposer utilize SD v1-5. More details on how we set up each method’s dual branch of PD and FD can be found in the supplementary material.
Evaluation Metrics. We adopt CLIP-T [16] to calculate prompt consistency (PC). To calculate identity fidelity (IF), we use MTCNN [51] and FaceNet  [37] to extract the embedding of the generated/reference faces and compute the cosine similarity. Following PhotoMaker, we adopt the face diversity (Face Div.) metric which calculates LPIPS [52] between facial areas. Lastly, following Face2Diffusion, we adopt PC $\times$ IF score, since it reflects the overall balance of prompt consistency and identity fidelity. We compute the harmonic mean (hMean) of PC and IF.
FreeCure Settings. We set $\omega$ in FASA to $2.0$, to ensure that attribute information from FD can be sufficiently integrated into PD. The $\gamma$ in APG is set to $0.5$ to keep the balance of identity fidelity and prompt consistency restoration. For attribute segmentation, we leverage BiSeNet [49] and Segment-Anything [21] for different facial attributes.
@@ 5.1 Main Results
Overall Performance Comparison. Table 1 and Figure 6 present a comparative analysis of our proposed method with the baselines, examining both quantitative and qualitative aspects. Generally, the baseline methods often fail to accurately reflect key facial attributes mentioned in the prompts. For instance, attributes with large areas (hair, sunglasses, etc.) often fail to align with the target prompts. FastComposer, Face-Diffuser, and PuLID consistently generate faces with identical expression. This observation reinforces our earlier discussion in Section 3. Conversely, our approach shows a remarkable ability to bolster absent or faint attributes, significantly enhancing prompt consistency for these baselines. Secondly, we notice that incorporating FreeCure leads to a slight decline in identity fidelity. This can be attributed to the increased facial diversity in the enhanced results. In essence, since baseline personalization models tend to produce faces closely resembling their references, they inherently score higher in identity fidelity. However, this does not align with the objective of creating new content that adheres to diverse prompts. Ultimately, in terms of the PC $\times$ IF metric, our method also shows considerable improvement over all baselines. Qualitatively, the enhancements are particularly evident for localized attributes like hair and accessories and abstract attributes including expressions. In conclusion, both quantitatively and qualitatively, FreeCure demonstrates a positive balance by enhancing prompt consistency while keeping the reduction in identity fidelity to a minimum. Additional comparisons and results on more non-celebrity references are available in Section 9 of the supplementary material.
@@
PuLID (PD) 25.16 58.23 42.12 35.14
PuLID + FreeCure 26.05 (+3.55%) 56.95 (-2.20%) 43.72(+3.80%) 35.74 (+1.74%)
Prompt Consistency with Multiple Attributes. Prompts involving multiple facial attributes pose a greater challenge to personalization’s prompt consistency but better reflect practical user needs. Table 2 illustrates FreeCure’s improvements on baselines with prompts including multiple facial attributes. As the number of attributes increases, the PC values of baselines tend to decrease. In contrast, FreeCure’s improvement in PC becomes more significant as the complexity of the prompt increases.
FastComposer (PD) 18.83 18.05 16.92
Face-Diffuser (PD) 21.68 20.98 18.02
Face2Diffusion (PD) 22.54 21.98 20.54
PhotoMaker (PD) 23.51 23.01 22.57
PhotoMaker + FreeCure 25.04 (+6.51%) 24.91 (+8.26%) 24.64 (+11.19%)
PuLID (PD) 25.56 25.12 24.43
PuLID + FreeCure 26.30 (+3.34%) 25.96 (+5.32%) 25.73 (+3.55%)
Figure 7: Demonstration of FreeCure’s non-interference manner: top row: personalization models’ outputs; middle row: intermediate enhanced results with FASA; bottom row: final results after APG enhancement.
@@ 5.2 Robustness Justification
Non-interference Enhancement. We validate that the subsequent enhancement processes of FreeCure do not impact attributes that have already been enhanced. Figure 7 visualizes the intermediate results at FreeCure’s different stages. When APG is used in the latter stages, attributes previously enhanced through FASA remain unaffected. This illustrates that FreeCure’s improvement across multiple attributes is highly robust and consistent.
Robust Performance with Different Initial Noises. We observe that, with identical initial noise, all baselines’ FD and PD processes generate faces with similar attribute locations. It is important to validate that FASA’s robust performance when faces generated by FD and PD have different locations and angles. We relax this condition by using different initial noises and regenerating the faces. Figure 8 shows that results under the two settings are comparable, which confirms that FASA can effectively enhance the generated results of PD, even when its FD counterpart produces faces with variable spatial structures.
@@
Visualization of FASA maps. Figure 9 visualizes the FASA map $A\in\mathbb{R}^{H\times(2\times W)}$, whose size is doubled according to Equation 5. For a given query point $q_{i}$ in $Q_{p}$, its corresponding scores $A_{i,:}$ are extracted from both $K_{p}$ and $K_{f}$. When $q_{i}$ falls into areas associated with target attributes (e.g., hair, sunglasses), the FASA map exhibits greater attention to information from $K_{f}$, corresponding to FD. Conversely, in regions unrelated to attributes, the FASA map retains its attention on PD. This visualization substantiates the role of FASA in transferring fine-grained attribute information, equivalent to detailed foundational knowledge, from FD to PD. Additionally, FASA preserves its original attention pattern in regions unrelated to attributes, thereby ensuring the performance of personalization models in maintaining identity fidelity. More detailed analyses are available in Section 10 of the supplementary material.
@@ 5.3 Ablation Study
Overall Analysis of Each Component. Table 3 presents the individual performance of FASA and APG in enhancing attributes across five baselines. Both components demonstrate positive effects compared to the baseline metrics in Table 1, with FASA showing more noticeable improvements. We attribute this to FASA’s ability to effectively handle attributes covering larger areas, such as hair and sunglasses, resulting in more observable enhancements.
FastComposer 20.91 41.94 27.91
19.03 42.84 26.35
Face-Diffuser 21.86 57.92 31.74
20.74 58.11 30.60
Face2Diffusion 22.73 39.30 28.80
21.98 39.47 28.24
PhotoMaker 24.76 50.62 33.25
23.61 50.86 32.25
PuLID 25.86 57.85 35.74
25.54 57.24 35.33
Table 3: Quantitative ablation analysis for FASA and APG’s independent effect on overall performance.
Figure 10: Ablation of scaling mask strategy.
@@
Scaling Mask Strategy. Figure 10 illustrates the evaluation of the scaling mask strategy implemented in FASA. Low scaling values critically hinder FASA’s capacity to effectively transfer attribute information to the personalization denoising. On the contrary, high scaling factors may negatively impact the overall quality of the generated faces. Generally, the optimal value of $\omega$ should be around $2.0$. More ablation studies of FASA are available in Section 10.3 of the supplementary material.
Figure 11: Ablation of APG with different intermediate timesteps.
@@
Effect of Inversion’s Intermediate Timesteps. Figure 11 demonstrates the effects of different starting timesteps in APG, represented by the parameter $\gamma$ in Section 4.2. To achieve a noticeable improvement in attributes while maintaining identity fidelity, an optimal $\gamma$ should be set as $0.5$.
6 Conclusion
@@
Facing the challenge that personalization models employing identity embeddings frequently struggle to preserve prompt consistency, we introduce FreeCure, a training-free framework that leverages the high prompt consistency inherent in foundation models to refine the output of personalization models, leading to a remarkable improvement in their prompt consistency. Additionally, FreeCure is able to enhance multiple attributes concurrently in a manner that is both complementary and non-disruptive. Our experiments validate the effectiveness of FreeCure on popular baselines particularly in scenarios with complex prompts that encompass multiple attributes.
Supplementary Material
7 Overview
The Supplementary Material consists of three sections, providing additional details and results to complement the main submission.
@@ 7 Overview
  • Section 8 provides implementation details, including details of setting up the dual denoising paradigm for the FD and PD processes, segmentation models as well as evaluation metrics.
@@
  • Section 9 provides additional results, including complementary FD outputs corresponding to the results included in Figure 6 of the main paper, additional experimental results on non-celebrity reference images and broader attribute enhancement achieved through the integration of advanced models, such as Segment-Anything.
@@
  • Section 10 provides a comprehensive analysis of FreeCure, including analysis for initial noise conditions, detailed visualization of FASA modules, additional ablation studies to determine the optimal placement of FASA within the U-Net, and additional investigations into identity embeddings.
8 More Implementation Details
8.1 Triggering Hidden Foundation Knowledge in Personalization Baselines

### v3

@@ 4.2 Asymmetric Prompt Guidance
Following attributes with clear location restored via FASA, to enhance more abstract attributes such as expressions, we introduce a simple yet effective method called asymmetric prompt guidance (APG). This strategy is based on the diffusion model’s inversion  song2020denoising . During the inversion phase, the model accepts only a template prompt (e.g. “a man") that does not include any target attributes. In the denoising process, we add the target attributes’ tokens back to this template prompt (e.g. “a man" $\rightarrow$ “a man laughing"). By leveraging the pretrained controllability of the foundation model, this approach enhances such attributes, resulting in the final refined image $I_{out}$. Throughout the denoising process, we use only pure textual prompts without identity embeddings, thereby avoiding their potential influence on the tokens related to the target attributes, as discussed in Sec.3. Furthermore, to better preserve the identity, we start the denoising process directly from an intermediate latent code $\hat{z_{\gamma T}},\gamma\in[0,1]$, where the high-level identity information has already been established.
@@ 5 Experiments
Evaluation Datasets. We collect an extensive dataset including $50$ identities, with $30$ derived from the CelebA-HQ CelebAMask-HQ and the other $20$ non-celebrity identities curated by our team. Each identity is represented by a single image, with a spectrum of facial characteristics. The prompt set consists of 20 prompts containing different facial attributes. For each (identity, prompt) pair, we produce 20 images. Detailed information can be found in Appendix.A.2.
Baselines. We evaluate FreeCure using several representative facial personalization methods: FastComposer xiao2024fastcomposer , Face-diffuser wang2024high , Face2Diffusion shiohara2024face2diffusion , InstantID wang2024instantid , PhotoMaker li2024photomaker , PuLID NEURIPS2024_409fcc9d , and InfiniteYou jiang2025infiniteyou . Among these, Face-diffuser, Face2Diffusion, and FastComposer are implemented using SDv1-5 rombach2022high , whereas InstantID, PhotoMaker, and PuLID are based on SD-XL podell2023sdxl . PuLID and InfiniteYou employ FLUX.1-dev bflflux as their foundation models.
Evaluation Metrics. We adopt CLIP-T hessel-etal-2021-clipscore to calculate prompt consistency (PC). To calculate identity fidelity (IF), we use MTCNN zhang2016joint and FaceNet  schroff2015facenet to extract the embedding of the generated/reference faces and compute the cosine similarity. Following PhotoMaker, we adopt the face diversity (Face Div.) metric which calculates LPIPS zhang2018unreasonable between facial areas. Lastly, following Face2Diffusion, we adopt PC $\times$ IF score, since it reflects the overall balance of prompt consistency and identity fidelity. We compute the harmonic mean (hMean) of PC and IF.
FreeCure Settings. We set $\omega$ in FASA to $2.0$, to ensure that attribute information from FD can be sufficiently integrated into PD. The $\gamma$ in APG is set to $0.5$ to maintain the balance between identity fidelity and prompt consistency. For attribute segmentation, we leverage BiSeNet yu2018bisenet and Segment-Anything kirillov2023segment for different facial attributes.
@@ 5.1 Main Results
PuLID (SDXL) 25.16 58.23 42.12 35.14
PuLID (SDXL) + FreeCure 26.05 (+3.55%) 56.95 (-2.20%) 43.72 (+3.80%) 35.74 (+1.74%)
PuLID (FLUX) 22.42 74.97 43.91 34.52
PuLID (FLUX) + FreeCure 24.78 (+10.53%) 72.61 (-3.15%) 46.09 (+4.96%) 36.95 (+7.04%)
InfiniteYou 23.77 79.71 44.28 36.62
InfiniteYou + FreeCure 25.25 (+6.23%) 77.13 (-3.24%) 46.82 (+5.74%) 38.05 (+3.90%)
Figure 5: Qualitative comparison with facial personalization baselines (including baselines built upon SDv1.5 and SDXL). Different attributes in prompts are highlighted in various colors. Comparison of corresponding FD outputs is provided in the Appendix.A.3.2.
Figure 6: Qualitative comparison with FLUX-based facial personalization baselines. Different attributes in prompts are highlighted in various colors.
Overall Performance Comparison. Table 1 and Fig. 5 & 6 present a comparative analysis of our proposed method with the baselines, examining both quantitative and qualitative aspects. It is easy to notice that baselines often fail to accurately reflect key facial attributes mentioned in the prompts. For instance, these baselines often generate faces with identical expression (row 2, 4 of Fig.5 and row 2 of Fig.6). For attributes with large areas (e.g., hair, sunglasses), baselines often cannot generate them harmoniously (row 3, 4 and 5 of Fig.5 and row 1 of Fig.6). Conversely, our approach shows a remarkable ability to enhance absent or faint attributes, significantly improving prompt consistency for these baselines. FreeCure can even tackle subtle facial attributes (e.g., eye color and earrings, see row 5, 6 of Fig.5). Notably, FreeCure achieves significant performance improvements across all foundation models’ personalization methods, demonstrating its strong generalizability. Secondly, we notice that FreeCure leads to a slight decline in identity fidelity, which can be attributed to positively improved facial diversity. Since baselines tend to produce faces closely resembling their references, they inherently score higher in identity fidelity. Ultimately, in terms of the PC $\times$ IF metric, our method also shows considerable improvement over all baselines. In conclusion, both quantitatively and qualitatively, FreeCure demonstrates a positive balance by enhancing prompt consistency while keeping the reduction in identity fidelity to a minimum. Additional comparisons and results on more references (including non-celebrities) are available in Appendix. A.3.1. We also provide corresponding FD outputs for Fig.5 to show the effectiveness of FreeCure to transfer correct attributes from the FD to PD process in Appendix. A.3.2
Prompt Consistency with Multiple Attributes. Prompts involving multiple facial attributes pose a greater challenge to personalization’s prompt consistency but better reflect practical user needs. Table 2 illustrates FreeCure’s improvements on baselines with prompts including multiple facial attributes. As the number of attributes increases, the PC values of baselines tend to decrease. In contrast, FreeCure’s improvement in PC becomes more significant as the complexity of the prompt increases. More analysis of FreeCure’s non-disruption manner in multiple prompt personalization is available in Appendix.A.4.1. In summary, through integration with FreeCure, personalization models can effectively address more complex and realistic prompt instructions.
Table 2: Quantitative comparison of prompt consistency with different number of attributes. After calculating the metrics for each method, we compute their mean values based on the corresponding foundation model type. For metrics of each baseline, please refer to Appendix.A.3.3.
Foundation Model PC (1 Attr.) $\uparrow$ PC (2 Attr.) $\uparrow$ PC (3 Attr.) $\uparrow$
SDv1.5 21.01 20.34 18.49
SDv1.5 + FreeCure 22.70 (+8.04%) 22.34 (+9.87%) 21.16 (+14.44%)
SDXL 23.83 23.31 22.65
SDXL + FreeCure 25.11 (+5.34%) 24.80 (+6.36%) 24.49 (+8.14%)
FLUX 24.15 22.64 21.88
FLUX + FreeCure 25.75 (+6.63%) 24.71 (+9.15%) 24.16 (+10.45%)
@@ 5.2 Robustness Justification
Robust Performance with Different Initial Noises. We observe that, with identical initial noise, all baselines’ FD and PD processes generate faces with similar attribute locations. It is important to validate that FASA’s robust performance without these condition. Thus, we relax this condition and regenerate faces with different initial noises. Fig.8 shows that results under the two settings are comparable, which confirms that FASA can effectively enhance the generated results of PD, even when its FD counterpart produces faces with variable spatial structures.
@@
Visualization of FASA maps. Fig.8 visualizes the FASA map $A\in\mathbb{R}^{H\times(2\times W)}$, whose size is doubled according to Eq.4. For a given query point $q_{i}$ in $Q_{p}$, its corresponding scores $A_{i,:}$ are extracted from both $K_{p}$ and $K_{f}$. When $q_{i}$ falls into areas associated with target attributes (e.g., hair, sunglasses), the FASA map exhibits greater attention to information from $K_{f}$, corresponding to FD. Conversely, in regions unrelated to attributes, the FASA map retains its attention on PD. This visualization substantiates the role of FASA in transferring fine-grained attribute information from FD to PD. Additionally, FASA preserves its original attention pattern in regions unrelated to attributes, thereby ensuring the performance of personalization models’ identity fidelity. More detailed visualization analyses of FASA are available in Appendix.A.4.6.
Table 3: Quantitative ablation analysis for FASA and APG’s independent effect. After calculating the metrics for each method, we compute their mean values based on the corresponding foundation model type. For metrics of each baseline, please refer to Appendix.A.4.4.
Foundation Model FASA APG PC(%) $\uparrow$ IF(%) $\uparrow$ PC $\times$ IF (hMean) $\uparrow$
SDv1.5 21.83 46.39 29.69
20.58 46.81 28.59
SDXL 24.55 56.97 34.31
23.97 56.87 33.73
FLUX 24.50 75.66 37.01
24.09 75.85 36.57
@@ 5.3 Ablation Study
Overall Analysis of Each Component. Table 3 presents the individual performance of FASA and APG in enhancing attributes across all baselines. Both components demonstrate positive effects compared to the baseline metrics in Table 1, with FASA showing more noticeable improvements. We attribute this to FASA’s ability to effectively handle attributes covering larger areas, such as hair and sunglasses, resulting in more observable enhancements.
(a) Ablation of FASA scaling mask strategy.
(b) Ablation of APG intermediate timesteps.
Figure 9: Ablation Studies.
@@
Scaling Mask Strategy. Figure 9(a) illustrates the evaluation of the scaling mask strategy implemented in FASA. Low scaling values critically hinder FASA’s capacity to effectively transfer attribute information to the personalization denoising. On the contrary, high scaling factors may negatively impact the overall quality of the generated faces. Generally, the optimal value of $\omega$ should be around $2.0$. More ablation studies of FASA are available in Appendix.A.4.5.
@@
Effect of Inversion’s Intermediate Timesteps. Figure 9(b) demonstrates the effects of different starting timesteps in APG, represented by the parameter $\gamma$ in Section 4.2. To achieve a noticeable improvement in attributes while maintaining identity fidelity, an optimal $\gamma$ should be set as $0.5$.
Hyperparameter Analysis. Our supplemental ablation analysis indicates that the hyperparameters $\omega$ and $\gamma$ consistently fall within specific ranges across different baselines: $\omega\in[1.8,2.4],\gamma\in[0.5,0.6]$. Baselines based on Stable Diffusion v1.5  shiohara2024face2diffusion ; xiao2024fastcomposer ; wang2024high require larger $\omega$ values (up to 2.4), while other methods  gu2024photoswap ; NEURIPS2024_409fcc9d ; jiang2025infiniteyou ; wang2024instantid perform well around $\omega=2.0$. Since results remain stable within these ranges, we adopt a unified setting ($\omega=2.0,\gamma=0.5$) throughout our experiments.
5.4 User Study
To further validate FreeCure’s superiority, we conducted an online user study with 30 participants. The study was designed as follows: for each baseline method, we randomly selected 10 samples. Each sample included a reference image, a text prompt, the baseline’s output, and the result refined by FreeCure. Participants were asked to evaluate the prompt consistency and identity fidelity of each output. As shown in Fig.10(a) and Fig.10(b), the results demonstrate a clear preference for FreeCure on prompt consistency and equal preference on identity fidelity, indicating that FreeCure’s potential negative impact on identity preservation is minimal.
(a) User preference on prompt consistency
(b) User preference on identity fidelity
Figure 10: User study of FreeCure. The preference ratio indicate that FreeCure can improve prompt consistency without undermining identity fidelity of different personalization models.
6 Limitation and Conclusion
Limitation. FreeCure is influenced by inherent bias and maximum achievable capabilities of personalization models, which is commonly observed in plug-and-play adaptation frameworks. Furthermore, it struggles with certain transparent objects, such as glass bottles. In some cases, we also observe slight attribute entanglement between FD and PD processes. Furthermore, FreeCure’s application to personalization models based on auto-regressive architectures could be further explored.
@@ 6 Limitation and Conclusion
Conclusion. Facing the challenge that personalization models employing identity embeddings frequently struggle to preserve prompt consistency, we introduce FreeCure, a training-free framework that leverages the high prompt consistency inherent in foundation models to refine the output of personalization models, leading to a remarkable improvement in their prompt consistency and minimal disruption to their original identity fidelity Our experiments validate the effectiveness of FreeCure on popular baselines from different foundation models, particularly in scenarios with complex prompts that encompass multiple attributes.
Acknowledgments
The research was supported in part by Theme-based Research Scheme (T45-205/21-N) from Hong Kong RGC, Generative AI Research, Development Centre from InnoHK, in part by the National Natural Science Foundation of China (Grant No. 62372480), in part by HKUST-MetaX Joint Lab Fund (No. METAX24EG01-D). We also thank Yubai Wei, Moheng Li, and Meixin Zhou for their assistance with the user study.
NeurIPS Paper Checklist
The checklist is designed to encourage best practices for responsible machine learning research, addressing issues of reproducibility, transparency, research ethics, and societal impact. Do not remove the checklist: The papers not including the checklist will be desk rejected. The checklist should follow the references and follow the (optional) supplemental material. The checklist does NOT count towards the page limit.
Please read the checklist guidelines carefully for information on how to answer these questions. For each question in the checklist:
  • You should answer [Yes] , [No] , or [N/A] .
  • [N/A] means either that the question is Not Applicable for that particular paper or the relevant information is Not Available.
  • Please provide a short (1–2 sentence) justification right after your answer (even for NA).
The checklist answers are an integral part of your paper submission. They are visible to the reviewers, area chairs, senior area chairs, and ethics reviewers. You will be asked to also include it (after eventual revisions) with the final version of your paper, and its final version will be published with the paper.
The reviewers of your paper will be asked to use the checklist as one of the factors in their evaluation. While "[Yes] " is generally preferable to "[No] ", it is perfectly acceptable to answer "[No] " provided a proper justification is given (e.g., "error bars are not reported because it would be too computationally expensive" or "we were unable to find the license for the dataset we used"). In general, answering "[No] " or "[N/A] " is not grounds for rejection. While the questions are phrased in a binary way, we acknowledge that the true answer is often more nuanced, so please just use your best judgment and write a justification to elaborate. All supporting evidence can appear either in the main paper or the supplemental material, provided in appendix. If you answer [Yes] to a question, in the justification please point to the section(s) where related material for the question can be found.
Appendix A Appendix
The Appendix consists of three sections:
  • Sec.A.1 includes an ethics statement that addresses the privacy of our evaluation datasets and our methods for mitigating potential risks.
@@ Appendix A Appendix
  • Sec.A.2 provides implementation details, including details of setting up the dual denoising paradigm for every baseline. Details of segmentation models, prompt datasets as well as evaluation metrics are also included.
@@
  • Sec.A.3 provides additional results of FreeCure in a boarder range of references, including non-celebrities. It also includes complementary FD outputs corresponding to the results of the main submission (Fig.5, 6), and broader attribute enhancement achieved through the integration of advanced models, such as Segment-Anything Model [28].
@@
  • Sec.A.4 provides a comprehensive analysis of FreeCure, including analysis for initial noise conditions, detailed visualization of FASA modules, more detailed ablation studies, runtime analysis and a template of the user study.
A.1 Ethics Statements
We confirm that our human face datasets, featuring both celebrities and non-celebrities, are sourced from public datasets or Google Images, limited to CC-licensed content. We have made extensive efforts to balance critical attributes like gender, race, and age to mitigate significant bias. Furthermore, we acknowledge that FreeCure could be misused for malicious purposes, such as creating impersonating identities. To mitigate these risks, we are committed to ethical governance and a controlled release strategy. For instance, we release the method only to accredited researchers under a license prohibiting misuse, with mandatory ethics training.
A.2 More Implementation Details
A.2.1 Triggering Hidden Foundation Knowledge in Personalization Baselines